# A mathematical theory of the critical point of ferromagnetic Ising systems

### v1

@@ Abstract
We develop a theory of the critical point of the ferromagnetic Ising model, whose basic objects are the ergodic (pure) states of the infinite system. It proves the existence of anomalous critical fluctuations, for dimension $\nu=2$ and, under a standard assumption, for $\nu=3$, for the model with nearest-neighbor interaction, in a way which is consistent with the probabilistic approach of Cassandro, Jona-Lasinio, and several others, reviewed in Jona-Lasinio’s article in Phys. Rep. 352,439 (2001). We propose to single out the state at the critical temperature $T_{c}$, among the ergodic thermal states associated to temperatures $0\leq T\leq T_{c}$, by a condition of non-summable clustering of the connected two-point function. The analogous condition on the connected (2r)- point functions, for $r\geq 2$ , together with a scaling hypothesis, proves that the (macroscopic) fluctuation state is quasi-free, also at the critical temperature. Other subjects treated include topics relating to universality, including spontaneous breaking of continuous symmetries and violations of universality in problems of energetic and dynamic stability.
@@ 1.1 Introduction and summary
As remarked by Glimm in his review of certain books and articles on statistical mechanics [Gli76], “statistical mechanics concerns infinite systems, and the relevant analysis is over infinite-dimensional spaces”. We shall use his illuminating introduction as a useful guide. Let $\Lambda$ a finite subset of $\mathbb{Z}^{\nu}$, with volume (identified with the number of sites) $|\Lambda|$, with $\nu\geq 2$ the dimension and $\eta(\Lambda)$ denote the set of nearest-neighbor pairs $x,y\in\Lambda$ with $\left|x-y\right|=1$. Consider the measures on $\mathbb{R}^{\left|\Lambda\right|}$
@@
are specified in (1.1.1) with the nearest-neighbors set $\eta(\Lambda)$ replaced by $\eta\left(\Lambda\cup\delta\Lambda\right)$. For high temperature $T$ in (1.1.1),(1.1.2) (“weak couplings” $J/T<<1$, $h/T<<1$),
@@
Note that the second notion of phase transition is independent of the first one, because $f(T,h)$ is independent of the b.c. [Thomp72].
@@
In the last half-century, this idea was considerably developed, perhaps most deeply, in the realm of the Ising model, through the probabilistic approach to the theory of the renormalization group (RG), championed by several people, among them M. Cassandro, G. Jona-Lasinio, G. Gallavotti, A. Martin-Löf, P. M. Bleher and Ya. G. Sinai ([CasJL78], [CasJL76], [GML75], [BleSi73], [BleSi75]), and reviewed by Jona-Lasinio in [JL01], to which we refer for numerous additional references. A basic result in this context was Bleher and Sinai’s treatment [BleSi73], [BleSi75] of the critical behavior of Dyson’s hierarchical model [Dy69], which was expected to describe qualitatively well the critical behavior of the Ising model with both short-range and long-range interactions. These seminal papers greatly stimulated further work on hierarchical models for classical spin systems, see, in particular, [MCG08] and references therein.
The measure $d\mu_{T}^{\Lambda}=d\mu_{T,h=0}^{\Lambda}$ generates a state $\omega_{T}^{\Lambda}$ over the (abelian) $\mathrm{C}^{\ast}$algebra of functions $f_{\Lambda}$ of the variables $\{\sigma_{x};\ x\in\Lambda\}$ by
@@
$\omega_{T}^{\Lambda}(f_{\Lambda})\geq 0\qquad\text{ if\qquad\ }f_{\Lambda}\geq 0\text{ and }\omega_{T}^{\Lambda}(\mathbf{1})=1$ (1.1.9)
@@
where the quotation marks above mean that $f$ is the limit of the $f_{\Lambda}$ in the sup norm ([RR67], Section 1). We may also view (this is the natural approach from a physical standpoint) the Ising model as an anisotropic limit of a quantum model: in this case (the forthcoming (1.2.1), one may speak of the “Ising universality class”, for which phase transitions with spontaneous symmetry breaking (s.s.b.) of a discrete symmetry, the spin-flip symmetry $\sigma_{x}\rightarrow-\sigma_{x}\forall x\in\mathbb{Z}^{\nu}$ (and its generalization in the quantum case) take place, see also Subsection 1.3.1. The finite-region algebras $\mathcal{A}_{\Lambda}$ are then the spin algebras generated by the Pauli matrices $\{\sigma_{x}^{1,2,3}\}$, and the “quasilocal algebra” $\mathcal{A}$ is the inductive limit of the $\mathcal{A}_{\Lambda}$ (i.e., the norm closure of the union of the $\mathcal{A}_{\Lambda}$, see Subsection 1.2. We adopt this view (which will be particularly useful for the dynamics in Subsection 3.2); restriction to the abelian algebra generated by the $\{\sigma_{x}^{3}\}$ coincides with the previous description.
The study of statistical mechanics from the point of view of the states of infinite systems began, in the Ising model, with the seminal paper of Lebowitz and Martin-Löf [LML72], followed by Bricmont, Fontaine and Landau [BFL79], for classical spin systems: they proved the unicity of the equilibrium state for $T>T_{c}$.
@@
The capital letter $W$ above refers to “Wightman” in order to emphasize the close connection with quantum field theory, see Subsection 1.4. Equation (1.1.11) is a general property of a special class of states, the extremal ergodic states $\omega^{\mathrm{e}}$. These states are identified with the pure phases [Wight79]: they cannot be written as a convex combination of two other states $\omega_{1}\neq\omega^{\mathrm{e}}$ and $\omega_{2}\neq\omega^{\mathrm{e}}$, i.e., the relation $\omega^{\mathrm{e}}=a\omega_{1}+(1-a)\omega_{2}\text{ with }0<a<1$ does not hold unless $\omega_{1}=\omega^{\mathrm{e}}$ and $\omega_{2}=\omega^{\mathrm{e}}$. We now restrict our discussion to the phase-transition region of the Ising model, the extremal ergodic states being $\omega_{1}^{\mathrm{e}}=\omega_{T}^{+}$ and $\omega_{2}^{\mathrm{e}}=\omega_{T}^{-}$; these states may be obtained as limits (as previously described) of the states $\omega_{T}^{\Lambda}$, as $\Lambda\nearrow\mathbb{Z}^{\nu}$, with $+$ or -- b.c., respectively, and are not invariant under the spin-flip symmetry, i.e., as stated before, there is s.s.b.. If $0\leq T<T_{c}$, the closed segment
@@
Thus, the critical point is singled out among the ergodic states by the condition of non-summable clustering (1.1.14).
@@
At the critical point we have [Thomp72]
@@
(provided they exist) need not coincide, see Subsection 1.3.3. Moreover, for instance, the limit (1.1.26) need not exist in general: for instance, taking free b.c. on $\Lambda$ on the r.h.s. of (1.1.26), for $T<T_{c}$, one might have $\omega_{\Lambda_{n},T}=a_{n}\omega_{\Lambda_{n},T}^{+}+(1-a_{n})\omega_{\Lambda_{n},T}^{-}$, with $10<a_{n}<$ 1, and $n=1,2,...$, leading to (possibly) several distinct subsequential limits along sequences of finite regions $\Lambda_{n},n=1.2,...$ of the r.h.s. of (1.1.26).
@@
Section 2 is devoted to the connected Wightman functions ($r\geq 2$), both for $0<T<T_{c}$ and for $T=T_{c}$. Subsection 2.1 describes the framework which was introduced by one of us (M.R.) in [MR02] and reviews the case $0<T<T_{c}$, both for $r=2$ and $r>2$. In Subsection 2.2 we restrict ourselves to $r=2$ and $T=T_{c}$, establishing the existence of anomalous fluctuations for $\nu=2$ and for $\nu=3$, under the standard assumption (2.3.24). In Subsection 2.3 we approach the question of the connected $r$-point function for $r\geq 4$ in the critical case. Under an assumption (Scaling Assumption C), which is the natural analogue of (2.3.16) for $r\geq 4$, we prove in Theorem 2.6 that the fluctuation state $\omega_{F}^{T}$ is quasi-free (corresponding to a Gaussian characteristic function).
In Section 3 we study several aspects of universality and its violation. Subsection 3.1 discusses the main features of universality and in Proposition 3.1 we show that Criterion A is invalid in the case of s.s.b. of a continuous symmetry. Using the Dyson model as a prototype the known universality of critical exponents such as $\beta$ and $\gamma$ for long-range interactions, is briefly reviewed in Subsection 3.2.1. In Subsection 3.2.2 we argue that certain properties of energetic and dynamical stability, which distinguish long-range interactions from short-range ones, are among those which violate universality. In order to prove this conjecture in an explicit model (the Dyson model), we show that the Cloitre function, recently introduced by Albert and Kiessling ([AlKi17], see also [Ki13]), allows an exact calculation of “dynamical critical exponents”, which are seen not to be universal, and, moreover, are proved to be totally unaccounted for by (rigorous) nonequilibrium mean-field theory.
Except for Theorem 1.1, Theorem 2.3 and the theorem in Appendix A, the theorems and propositions in this paper are new, but it seemed more appropriate to write the paper as a review for reasons of clarity.
@@
Appendix B provides (we hope) a comprehensive review of the beautiful work of Newman [Ne79]. Remark B.2 explains in some detail the differences between his framework and ours. Newman’s results, perhaps surprisingly, turn out to complement ours in an essential way (see Remark 1.4 and Remark B.2).
Since we shall base our whole discussion on the states of infinite systems, which are the generalization of (1.1.10) to the whole “quasilocal” algebra $\mathcal{A}$, we need to make an interlude in the next subsection in order to establish some of the fundamental notions.
1.2 Mathematical interlude on states of infinite systems
@@
Above, $\left|x\right|\equiv\left(\sum_{i=1}^{\nu}x_{i}^{2}\right)^{1/2}$, $\nu$ denotes the dimension, and, as asserted in the introduction, the external field will be set equal to zero in the present paper. Equation (1.2.4) is the important requirement of thermodynamic stability, necessary for the existence of the thermodynamic limit of the energy per unit volume. Above, $\delta\in(-1,1)$ is the anisotropy ([BRo297], p. 320), and we assume
@@
which is appropriate to describe ferromagnetism. Above, too, $\sigma_{x}^{i}$, $i=1,2,3$, are the Pauli matrices at the site $x$, and $H_{\Lambda}$ acts on the Hilbert space $\mathcal{H}_{\Lambda}=\otimes_{x\in\Lambda}\mathbb{C}_{x}^{2}$, and $\vec{S}_{x}$ is short for $\mathbf{\ 1}\otimes\cdots\otimes\vec{S}_{x}\otimes\cdots\otimes\mathbf{1}$. We define the algebra of (bounded) operators acting on the Hilbert space $\mathcal{H}_{\Lambda}$: $\mathcal{A}(\Lambda)=B(\mathcal{H}_{\Lambda})$. These local algebras satisfy: 1.) Causality: $[\mathcal{A}(B),\mathcal{A}(C)]=0$ if $B\cap C=\phi$ and 2.) Isotony: $B\subset C\Rightarrow\mathcal{A}(B)\subset\mathcal{A}(C)$. $\mathcal{A}_{L}=\cup_{B}\mathcal{A}(B)$ is termed the local algebra; its closure with respect to the norm, the quasilocal algebra (observables which are, to arbitrary accuracy, approximated by observables attached to a finite region). The norm is $B(\mathcal{H}_{\Lambda})\ni A\longmapsto\left\|A\right\|=\sup_{\left\|\Psi\right\|\leq 1}\left\|A\Psi\right\|$, $\Psi\in\mathcal{H}_{\Lambda}$.
For a finite quantum spin system let $\rho_{\Lambda}$ denote a density matrix on $\mathcal{H}_{\Lambda}$, i.e., a non-negative Hermitian matrix of unit trace on $\mathcal{H}_{\Lambda}$. We may view $\rho_{\Lambda}$ as a state $\omega_{\Lambda}$ on $\mathcal{A}(\Lambda)$ – a positive, normed linear functional on $\mathcal{A}(\Lambda)$: $\omega_{\Lambda}(A)=\text{tr}_{\mathcal{H}_{\Lambda}}(\rho_{\Lambda}A)\text{ for }A\in\mathcal{A}(\Lambda)$ (positive means $\omega_{\Lambda}(A^{{\dagger}}A)\geq 0$, normed $\omega_{\Lambda}(\mathbf{1})=1$.)
This notion of state generalizes to systems with infinite number of degrees of freedom $\omega(A)=\lim_{\Lambda\nearrow\infty}\omega_{\Lambda}(A)$, at first for $A\in\mathcal{A}_{L}$ and then to $\mathcal{A}$.
@@
An automorphism $\tau(A)$ of $\mathcal{A}$ is basic concept: it is a one-to-one mapping of $\mathcal{A}$ onto $\mathcal{A}$ which preserves the algebraic structure. For quantum spin systems, we speak of  time-translation automorphisms $\tau_{t}$, defined as the limit in the operator norm of the time-evolutes $\exp(iH_{\Lambda}t)A\exp(-iH_{\Lambda}t)$ as the region $\Lambda\nearrow\infty$, and one speaks of a C*-dynamical system $(\mathcal{A},\tau_{t})$. The space-translation automorphisms $\tau_{x}$ are defined on $\mathcal{A}_{L}$, and subsequently by extension, by $(\tau_{x}A)(y)=A(x+y)\text{ for }x\text{ in a bounded region}$. The letter $\nu$ will always denote the spatial dimension.
Each state $\omega$ defines a representation $\Pi_{\omega}$ of $\mathcal{A}$ as bounded operators on a Hilbert space $\mathcal{H}_{\omega}$ with cyclic vector $\Omega_{\omega}$ (i.e., $\Pi_{\omega}(\mathcal{A})\Omega_{\omega}$ is dense in $\mathcal{H}_{\omega}$), such that $\omega(A)=(\Omega_{\omega},\Pi_{\omega}(A)\Omega_{\omega})$ (the GNS construction: see [BRo186], p.153). The strong closure of $\Pi_{\omega}(\mathcal{A})$ is a von Neumann algebra, with commutant $\Pi_{\omega}(\mathcal{A})^{{}^{\prime}}$, which is the set of bounded operators on $\mathcal{H}_{\omega}$ which commute with all $\Pi_{\omega}(\mathcal{A})$, and the center is defined by $Z_{\omega}=\Pi_{\omega}(\mathcal{A})\cap\Pi_{\omega}(\mathcal{A})^{{}^{\prime}}$.
@@
As remarked in the previous subsection, an extremal invariant or ergodic state is a state which cannot be written as a proper convex combination of two distinct states $\omega_{1}$ and $\omega_{2}$, i.e., the following does not hold:
@@
A factor or primary state is defined by the condition that the center
$Z_{\omega}=\{\lambda\mathbf{1}\}$ (1.2.9)
with $\lambda\in\mathbb{C}$.
For quantum spin systems the center $Z_{\omega}$ coincides with the so called algebra at infinity $\zeta_{\omega}^{\perp}$, which corresponds to operations which can be made outside any bounded set. The strong limit slim below is the limit when applied to any vector in the representation space. The following theorem is fundamental ([KaRo66], see also [LvH78] for an elementary proof):
@@
with $A=\sigma^{3}$. If $\eta_{\omega_{+}}(\sigma^{3})=a_{+}=1$, and $\eta_{\omega_{-}}(\sigma^{3})=-1$, the states $\omega_{\pm}$ are macroscopically different, i.e., differ from one another by flipping an infinite number of spins. For a comprehensive discussion, see [Sew86], Section 2.3.
Finally, a word about equilibrium states is necessary (for much more, see [Hug72]). There is an algebra $\tilde{\mathcal{A}}$ of analytic elements which is dense in $\mathcal{A}$, i.e., such that the function $\tau_{z}(A)$ is an entire function of the complex variable $z$ with values in $\mathcal{A}$, and reduces to the time-translation automorphism of $\mathcal{A}$ when $z$ is real (Lemmas 4.7 and 4.8 of [Hug72]). We say that a state $\omega$ is a thermal equilibrium state of the system at inverse temperature $\beta$, or, equivalently, a KMS state with respect to $\tau_{t}$ if
$\omega(\tau_{t}(A)B)=\omega(B\tau_{t+i\beta}(A))\qquad\forall B\in{\mathcal{A}}\mbox{ and }A\in\tilde{\mathcal{A}}$ (1.2.13)

### v2

@@ Abstract
We develop a theory of the critical point of the ferromagnetic Ising model, whose basic objects are the ergodic (pure) states of the infinite system. It proves the existence of anomalous critical fluctuations, for dimension $\nu=2$ and, under a standard assumption, for $\nu=3$, for the model with nearest-neighbor interaction, in a way which is consistent with the probabilistic approach of Cassandro, Jona-Lasinio, and several others, reviewed in Jona-Lasinio’s article in Phys. Rep. 352,439 (2001). We propose to single out the state at the critical temperature $T_{c}$, among the ergodic thermal states associated to temperatures $0\leq T\leq T_{c}$, by a condition of non-summable clustering of the connected two-point function. The analogous condition on the connected (2r)- point functions, for $r\geq 2$ , together with a scaling hypothesis, natural within our framework, proves that the (macroscopic) fluctuation state is quasi-free, after a proper rescaling, also at the critical temperature, in agreement with a theorem by Cassandro and Jona-Lasinio, whose proof is, however, shown to be incomplete. Other subjects treated include topics relating to universality, including spontaneous breaking of continuous symmetries and violations of universality in problems of energetic and dynamic stability.
Keywords: critical point; ferromagnetic Ising model; phase transition; criterion of criticality; fluctuation states; universality.
@@ 1.1 Introduction and summary
As remarked by Glimm in his review of certain books and articles on statistical mechanics [Gli76], “statistical mechanics concerns infinite systems, and the relevant analysis is over infinite-dimensional spaces”. We shall use his illuminating introdution as a useful guide. Let $\Lambda$ a finite subset of $\mathbb{Z}^{\nu}$, with volume (identified with the number of sites) $|\Lambda|$, with $\nu$ the dimension and $\eta(\Lambda)$ denote the set of nearest-neighbor pairs $x,y\in\Lambda$ with $\left|x-y\right|=1$. Consider the measures on $\mathbb{R}^{\left|\Lambda\right|}$
@@
are specified in (1.1.1) with the nearest-neighbors set $\eta(\Lambda)$ replaced by $\eta\left(\Lambda\cup\delta\Lambda\right)$. For high temperature $T$ in (1.1.1),(1.1.2) (“weak couplings” $J/T<<1$), $h=0$, or any value of $J$ if $h\neq 0$),
@@
Note that the second notion of phase transition is different from the first one, because $f(T,h)$ is independent of the b.c. [Thomp72].
Note also that in (1.1.7), the condition $m_{s}>0$ characterizes a first-order transition, which occurs for $0\leq T<T_{c}$, but, at the critical point, $m_{s}=m_{s}(T)\to 0$ as $T\to T_{c}^{-}$ (also as $T\to T_{c}^{+}$) continuously, see the forthcoming Remark 1.4 and the reference given there. Near $T=T_{c}$, we have the forthcoming (1.1.17), i.e., the susceptibility diverges, and therefore the critical point is a second-order point. This is further confirmed by looking at the manifold of points $(T,h)$, the specific heat diverges at $T=T_{c}$ when $h=0$.
@@
In the last half-century, this idea was considerably developed, perhaps most deeply, in the realm of the Ising model, through the probabilistic approach to the theory of the renormalization group (RG), championned by several people, among them M. Cassandro, G. Jona-Lasinio, G. Gallavotti, A. Martin-Löf, P. M. Bleher and Ya. G. Sinai ([CasJL78], [CasJL76], [GML75], [BleSi73], [BleSi75]), and reviewed by Jona-Lasinio in [JL01], to which we refer for numerous additional references. A basic result in this context was Bleher and Sinai’s treatment [BleSi73], [BleSi75] of the critical behavior of Dyson’s hierarchical model [Dy69], which was expected to describe qualitatively well the critical behavior of the Ising model with both short-range and long-range interactions. These seminal papers greatly stimulated further work on hierachical models for classical spin systems, see, in particular, [MCG08] and references therein.
The measure $d\mu_{T}^{\Lambda}=d\mu_{T,h=0}^{\Lambda}$ generates a state $\omega_{T}^{\Lambda}$ over the (abelian) $\mathrm{C}^{\ast}$-algebra of functions $f_{\Lambda}$ of the variables $\{\sigma_{x};\ x\in\Lambda\}$ by
@@
$\omega_{T}^{\Lambda}(f_{\Lambda})\geq 0\qquad\text{ if }\qquad f_{\Lambda}\geq 0\quad\text{ and }\quad\omega_{T}^{\Lambda}(\mathbf{1})=1$ (1.1.9)
@@
where the quotation marks above mean that $f$ is the limit of $f_{\Lambda}$ in the sup norm ([RR67], Section 1). We may also view (this is the natural approach from a physical standpoint) the Ising model as an anisotropic limit of a quantum model: in this case (the forthcoming (1.2.1), one may speak of the “Ising universality class”, for which phase transitions with spontaneous symmetry breaking (s.s.b.) of a discrete symmetry, the spin-flip symmetry $\sigma_{x}\rightarrow-\sigma_{x},~{}\forall x\in\mathbb{Z}^{\nu}$ (and its generalization in the quantum case) take place, see also Subection 1.3.1. The finite-region algebras $\mathcal{A}_{\Lambda}$ are then the spin algebras generated by the Pauli matrices $\{\sigma_{x}^{1,2,3}\}$, and the “quasilocal algebra” $\mathcal{A}$ is the inductive limit of the $\mathcal{A}_{\Lambda}$ (i.e., the norm closure of the union of the $\mathcal{A}_{\Lambda}$, see Subsection 1.2. We adopt this view (which will be particularly useful for the dynamics in Subsection 3.2; restriction to the abelian algebra generated by the $\{\sigma_{x}^{3}\}$ coincides with the previous description.
The study of statistical mechanics from the point of view of the states of infinite systems in both classical and quantum systems has a long history, beginning with the paper of Haag, Hugenholtz and Winnink [HHW67], and the papers of Lanford and Ruelle [LRu69] (see also Ruelle’s book [Rue69]), and going over to the results of uniqueness of the equilibrium state of the Ising model for $T>T_{c}$ due to Dobrushin [D72] and Lebowitz and Martin-Löf [LML72]. The books by Simon [Sim93] and Israel [Isr79] are good sources from the point of view of treating classical and quantum models in parallel.
@@
The capital letter $W$ above refers to “Wightman” in order to emphasize the close connection with quantum field theory, see Subsection 1.4. Other terminologies in common use are “cumulants”, “truncated correlatin functions”, and “Ursell functions”. Equation (1.1.11) is a general property of a special class of states, namely those ergodic states $\omega^{\mathrm{e}}$, which are also in the subset of the set of extremal states which are invariant under a group (see, e.g. [Rue69], Definition 6.3.1). In our case, the group is that of lattice translations on $\mathbf{Z}^{\nu}$, and the extremal states are identified with the pure phases [Wight79]: they cannot be written as a convex combination of two other states $\omega_{1}\neq\omega^{\mathrm{e}}$ and $\omega_{2}\neq\omega^{\mathrm{e}}$, i.e., the relation $\omega^{\mathrm{e}}=a\omega_{1}+(1-a)\omega_{2}\text{ with }0<a<1$ does not hold unless $\omega_{1}=\omega^{\mathrm{e}}$ and $\omega_{2}=\omega^{\mathrm{e}}$. We now restrict our discussion to the phase-transition region of the Ising model, the extremal ergodic states being $\omega_{1}^{\mathrm{e}}=\omega_{T}^{+}$ and $\omega_{2}^{\mathrm{e}}=\omega_{T}^{-}$; these states may be obtained as limits (as previously described) of the states $\omega_{T}^{\Lambda}$, as $\Lambda\nearrow\mathbb{Z}^{\nu}$, with $+$ or -- b.c., respectively, and are not invariant under the spin-flip symmetry, i.e., as stated before, there is s.s.b.. If $0\leq T<T_{c}$, the closed segment
@@
Remark 1.1.
It should be emphasized that the concepts of extremal and ergodic states may be distinct, as, for instance, in the Ising antiferromagnet, for which there are two invariance groups, corresponding to the translations within each sublattice. In this paper we shall, however, restrict ourselves to the Ising universality class, for which the two concepts coincide, and use the words “extremal states” and “ergodic states” interchangeably.
@@
Thus, the critical point is singled out among the ergodic states by the condition of non-summable clustering (1.1.14). In the one-dimensional $1/r^{2}$ model Imbrie and Newman ([IN88], see also [DHUM92]) have proved, however, that there is an intermediate regime $T_{0}<T<T_{c}$ with positive magnetization but nonsummable connected two-point functions.
Remark 1.2.
Although not for Ising models, it is possible that there is a first-order transition, as happens for n.n. high-q Potts models. At the “critical” temperature separating the low-T and high-T regime, there coexist q extremal (ergodic) ordered states with one ergodic (extremal) disordered state, all $q+1$ states with exponentially decaying correlations. We thank the referee for this remark.
@@
Near the critical point we have [Thomp72]
@@
(provided they exist) need not coincide, see Subsection 1.3.3. Moreover, for instance, the limit (1.1.26) need not exist in general: for instance, taking random b.c. on $\Lambda$ on the r.h.s. of (1.1.26), for $T<T_{c}$, one may have several distinct subsequential limits along sequences of finite regions $\Lambda_{n},n=1.2,...$ of the r.h.s. of (1.1.26), see [vES02].
@@
Section 2 is devoted to the connected Wightman functions ($r\geq 2$), both for $0<T<T_{c}$ and for $T=T_{c}$. Subsection 2.1 describes the framework which was introduced by one of us (M.R.) in [MR02] and reviews the case $0<T<T_{c}$, both for $r=2$ and $r>2$. In Subsection 2.2 we restrict ourselves to $r=2$ and $T=T_{c}$, establishing the existence of anomalous fluctuations for $\nu=2$ and for $\nu=3$, under the standard assumption (2.3.24).
In Subsection 2.3 we approach the question of the connected $r$-point function for $r\geq 4$ in the critical case. Under an assumption (Scaling Assumption C), which is the natural analogue of (2.3.16) for $r\geq 4$, we prove in Theorem 2.6 that the fluctuation state $\omega_{F}^{T}$ is quasi-free (corresponding to a Gaussian characteristic function). This conjecture seems to be widespread among several mathematical physicists (private communications by S. Molchanov and G. Slade to one of us (D.H.U.M.)), but the sole theorem we were able to found which claims to prove it was Theorem 6.2 in the paper by Cassandro and Jona-Lasinio [CasJL76]. Their proof is, however, incomplete (see also  1.7 and  1.8 ,in particular concerning the possible connection with one of the peculiarities pointed out by Griffiths and Pearce, and the important work of Wehrl).
In Section 3 we study several aspects of universality and its violation. Using the Dyson model as a prototype the known universality of critical exponents such as $\beta$ and $\gamma$ for long-range interactions, is briefly reviewed in Subsection 3.1. In Subsection 3.2 we argue that certain properties of energetic and dynamical stability, which distinguish long-range interactions from short-range ones, are among those which violate universality. In order to prove this conjecture in an explicit model (the Dyson model), we show that the Cloitre function, recently introduced by Albert and Kiessling ([AlKi17], see also [Ki13]), allows an exact calculation of “dynamical critical exponents”, which are seen not to be universal, and, moreover, are proved to be totally unaccounted for by (rigorous) nonequilibrium mean-field theory.
@@
Appendix B provides (we hope) a comprehensive review of the beautiful work of Newman [Ne79], which is complementary to ours, in that it concerns the upper critical dimension.
Since we shall base our discussion on the states of infinite systems, which are the generalization of (1.1.10) to the whole “quasilocal” algebra $\mathcal{A}$, we need to make an interlude in the next subsection in order to establish some of the fundamental notions. The notions need not be very extended, because we shall restrict ourselves to the Ising universality class (the forthcoming generalized Ising model in the case of nearest-neighbor interactions, the Dyson model in the case of long-range interactions).
1.2 Brief mathematical interlude on states of infinite systems
@@
Above, $\left|x\right|\equiv\left(\sum_{i=1}^{\nu}x_{i}^{2}\right)^{1/2}$, $\nu$ denotes the dimension, and, as asserted in the introduction, the external field will be set equal to zero in the present paper. Equation (1.2.4) is the important requirement of thermodynamic stability, necessary for the existence of the thermodynamic limit of the energy per unit volume. Above, $\delta\in(-1,1)$ and we assume
@@
which is appropriate to describe ferromagnetism. Above, too, $\sigma_{x}^{i}$, $i=1,2,3$, are the Pauli matrices at the site $x$, and $H_{\Lambda}$ acts on the Hilbert space ${\cal H}_{\Lambda}=\otimes_{x\in\Lambda}\mathbb{C}_{x}^{2}$, and $\vec{S}_{x}$ is short for $\mathbf{\ 1}\otimes\cdots\otimes\vec{S}_{x}\otimes\cdots\otimes\mathbf{1}$. We define the algebra of (bounded) operators acting on the Hilbert space ${\cal H}_{\Lambda}$: ${\cal A}(\Lambda)=B({\cal H}_{\Lambda})$. These local algebras satisfy: 1.) Causality: $[{\cal A}(B),{\cal A}(C)]=0$ if $B\cap C=\phi$ and 2.) Isotony: $B\subset C\Rightarrow{\cal A}(B)\subset{\cal A}(C)$. ${\cal A}_{L}=\cup_{B}{\cal A}(B)$ is termed the local algebra; its closure with respect to the norm, the quasilocal algebra (observables which are, to arbitrary accuracy, approximated by observables attached to a finite region). The norm is $B({\cal H}_{\Lambda})\ni A\longmapsto\left\|A\right\|=\sup_{\left\|\Psi\right\|\leq 1}\left\|A\Psi\right\|$, $\Psi\in{\cal H}_{\Lambda}$.
For a finite quantum spin system let $\rho_{\Lambda}$ denote a density matrix on ${\cal H}_{\Lambda}$, i.e., a non-negative Hermitian matrix of unit trace on ${\cal H}_{\Lambda}$. We may view $\rho_{\Lambda}$ as a state $\omega_{\Lambda}$ on $\cal{A}(\Lambda)$ – a positive, normed linear functional on ${\cal A}(\Lambda)$: $\omega_{\Lambda}(A)=\text{tr}_{{\cal H}_{\Lambda}}(\rho_{\Lambda}A)\text{ for }A\in{\cal A}(\Lambda)$ (positive means $\omega_{\Lambda}(A^{{\dagger}}A)\geq 0$, normed $\omega_{\Lambda}(\mathbf{1})=1$.)
This notion of state generalizes to systems with infinite number of degrees of freedom $\omega(A)=\lim_{\Lambda\nearrow\infty}\omega_{\Lambda}(A)$, at first for $A\in{\cal A}_{L}$ and then to ${\cal A}$.
@@
As remarked in the previous subsection, an extremal invariant or ergodic state (see, however, Remark 1.1) is a state which cannot be written as a proper convex combination of two distinct states $\omega_{1}$ and $\omega_{2}$, i.e., the following does not hold:
@@
The strong limit s-lim below is the limit when applied to any vector in the representation space. The following theorem is fundamental ([KaRo66], see also [LvH78] for an elementary proof):
@@
with $A=\sigma^{3}$. If $\eta_{\omega_{+}}(\sigma^{3})=a_{+}>0$, and $\eta_{\omega_{-}}(\sigma^{3})=a_{-}<0$, the states $\omega_{\pm}$ are macroscopically different, i.e., differ from one another by flipping an infinite number of spins. For a comprehensive discussion, see [Sew86], Section 2.3.
Finally, a word about equilibrium states is necessary (for much more, see [Hug72]). There is an algebra $\tilde{{\cal A}}$ of analytic elements which is dense in ${\cal A}$, i.e., such that the function $\tau_{z}(A)$ is an entire function of the complex variable $z$ with values in ${\cal A}$, and reduces to the time-translation automorphism of ${\cal A}$ when $z$ is real (Lemmas 4.7 and 4.8 of [Hug72]). We say that a state $\omega$ is a thermal equilibrium state of the system at inverse temperature $\beta$, or, equivalently, a KMS state with respect to $\tau_{t}$ if
$\omega(\tau_{t}(A)B)=\omega(B\tau_{t+i\beta}(A))\qquad\forall B\in{\cal A}\text{ \quad and \quad}A\in\tilde{{\cal A}}$ (1.2.12)